what is x??????
14 - 6x = 20 + 4x
the answer is 3 / -5
i want to know how the answer is that
so some one can u please help me
im trying to study and practise maths cause tomorrow is test
so i need to practise please help

Answer by computergeek008(7)

14 - 6x = 20 + 4x
Move all the terms with x on the left side, and the other terms on the right side. Then isolate x.

Subtract 4x from both sides.
14 - 10x = 20

Subtract 14 from both sides.
-10x = 6

Divide both sides by -10 to isolate the x
x = -6/10

Reduce the fraction
x = -3/5

Note: -3/5 is the same as 3/-5